# Anchored Learning for On-the-Fly Adaptation — Extended Technical Report

### v2

@@ IV-B Catastrophic Forgetting in Reacher
In Fig. 4, we showcase the severity of catastrophic forgetting when fine-tuning agents naively. This issue is especially grave on the reacher task since the target domain has exactly the same dynamics as the source domain, only differing in setpoint locations such that the ones in the target domain are a strict subset of the ones in the source. The unstable behavior of the naively fine-tuned agent is a result of the agent overfitting to the target domain and forgetting how to solve the task in the source domain.
IV-C Preventing Catastrophic Forgetting on Gymnasium
@@ V Anchor Critics in Reality
    Leveraging SwaNNL, we evaluate how live-adapting with and without anchors affects safety, smoothness, and performance on the real drone system in Sections VI-B and VI-C. The novel infrastructure offered by the SwaNNL enabled us to develop and test practical ways to adapt agents while a quadrotor is in flight, demonstrating how anchors make live adaptation safer on real drones.
@@ V-A Engineering SwaNN Lake for Live NN updates
In this work, we investigate strategies for RL-based live control adaptation on real robots. As such, two important sub-goals are formulated as follows. First, we needed to allow BetaFlight [34]-compatible firmware to update its NN controller(s) without interrupting the control loop. As BetaFlight is a standard across various robotic platforms, this allows our framework to support a large range of robots, such as blimps, boats, helicopters, racing drones, and more. Second, we needed to be able adapt the agents live without crashing. This necessitated preventing catastrophic forgetting.
We thus decided to evolve Neuroflight [10, 11], initially developed by Koch et al. for allowing the use of NNs for high-performance quadrotor control. Neuroflight extends the open-source Betaflight [34] flight-control firmware stack allowing NN models to be compiled and embedded into the control firmware. We evolved the Neuroflight framework such that modifications to the employed NN controller(s) can be carried out without interrupting the control loop. To do so, we re-designed the firmware so that we could swap out binarized NN graphs with minimal overhead rather than having them directly compiled into the firmware. In particular, it is possible to both (i) update the NN controller weights and (ii) swap-in graphs with different NN architectures. Our novel firmware stack outlined in Fig. 6, is open-sourced [30]. Our architecture allows inference to be performed onboard the micro-controller, but training onboard is prohibitively expensive, given compute and storage limitations of low-power embedded platforms. We thus introduce a point-to-point communication channel with an external ground station, to mitigate this issue. Most importantly, the ground station can be used as a training platform to adapt controllers based on data captured in flight. SwaNNFlight further integrates with wireless transceiver modules to (i) wirelessly transmit flight data and (ii) accept new NN models to switch to mid-flight.
@@ VI TRAINING FLIGHT CONTROLLERS WITH SwaNNL
Transitioning to TF2 introduced a problem for the training pipeline. Prior works [36, 35, 11] train flight-control agents with the OpenAI Baselines [37] implementation of PPO [19]. While reproducible with the software packages and versions reported by the authors, the performance observed with PPO agents in TF1 was not achievable by any of several tested TF2 implementations of PPO, including TF-agents [38], stable-baselines [39] etc., despite using identical network architectures and performing a full hyperparameter search. Algorithms like PPO and TRPO [18] are known to be sensitive to (and even brittle under) implementation specifics [40, 41, 42], prompting a search for alternatives.
Viable flight-control policies were trainable using a TF2 translation of OpenAI’s Spinning Up [43] PyTorch implementation of DDPG [31]. While vanilla DDPG presented with similar oscillatory behavior identified in the work on CAPS  [36], introducing CAPS regularization to DDPG enabled the training of viable agents in simulation. We found, however, that the performance of DDPG on the control problem at hand exhibited relatively high variance: not all the trained agents were safely flyable. This prompted us to explore alternative formulations which improve upon CAPS defined in the next section.
@@ VI-B Catastrophic Forgetting and Safety Concerns in Adaptation without Anchors
Before testing adaptation in live, unconstrained flight, we first conducted a series of tests in a controlled lab environment where the drone was attached to a light tether to restrain it in the event of a catastrophic loss of control, without being a hindrance to flight. These tests consisted of fine-tuning a well-trained controller from simulation (the source domain) on the real platform (the target domain) while giving it typically small control targets ($<50$ deg/s) and occasionally requesting control in excess of that ($>100$ deg/s). Without anchors preserving agents’ performance on larger control inputs—like those experienced during the simulated training—the live agents were observed to forget how to maintain tracking when receiving large inputs, though sometimes regaining control, and often losing it again, as seen in Fig. 7. This also results in a sudden loss of control where errors would quickly snowball and motor actuation rises sharply, risking damage to the drone. Fig. 8 demonstrates how, when commanding attitudes not seen for a while during live-flight adaptation, agents could very quickly lose control (observed from MAEs increasing quickly and exponentially) – which we attributed to overfitting over the relative homogeneity of the live-flight training data akin to the reacher environment we tested. Anchors help avoid this problem.
@@ VI-C Live Adaptation During Unconstrained Flight
We use two key performance metrics in policy evaluation. (1) Tracking accuracy is measured using the Mean Absolute Error (MAE). (2) Smoothness is computed by taking the Fast Fourier Transform (FFT) of the actuation commands, and then computing $Sm$ as defined by Equation 4 in [36] (lower is better). Shown in Fig. 10 is a comparison of control signals before and after 15 adaptation steps. Motor usage is reduced significantly, resulting in improved control smoothness and almost a 50% reduction in power consumption, with comparable MAEs to agents trained in simulation. This is further demonstrated in Table II, which examines how MAE, smoothness, and current draw change through adaptation.
VII Multiplicative Composition of Optimization Objectives
Multiplicative composition of multi-objective rewards through a geometric mean, with each reward in [0,1], has been shown to improve training performance and reduce policy variance compared to linear composition [35]. This approach, analogous to a smooth AND operator, incentivizes policies to satisfy all defined objectives. We apply this concept to our optimization objectives and regularization terms, improving upon the linear composition used in CAPS [36].
For an RL policy $\pi_{\theta}$, the DDPG policy optimization criteria under our multiplicative version of the CAPS regularization scheme is:
$J_{\pi_{\theta}}=\sqrt[4]{Q_{\pi_{\theta}}\left(\frac{w_{T}}{w_{T}+L_{T}}\right)\left(\frac{w_{S}}{w_{S}+L_{S}}\right)\left(\frac{w_{A}}{w_{A}+\pi_{\theta}^{:-1}}\right)}.$ (12)
where: $L_{T}$ and $L_{S}$ are temporal and spatial smoothness regularization terms, $Q_{\pi}$ is the critic estimating Q-values of state-action pairs, $\pi_{\theta}^{:-1}$ represents the policy network output prior to the final $\tanh$ activation, $w_{T}$, $w_{S}$, and $w_{A}$ are penalty thresholds for $L_{T}$, $L_{S}$ and $\pi_{\theta}^{:-1}$ respectively.
@@ VII Multiplicative Composition of Optimization Objectives
[…]
VII-A Ablation Study on Multiplicative Composition
To motivate the use of multiplicative composition in our optimization criteria, we trained controllers in simulation using CAPS-like [36] linear loss composition against multiplicative loss composition, shown in Equation 12. We conducted extensive hyperparameter search with different weighting parameters, both for the CAPS regularization terms as well as the training rewards. Even with our best set of hyperparameters, however, variance in trained performance remained high. We show in Fig. 9 that we achieved higher performance with less tuning and significantly reduced variance using multiplicative composition. The results presented in the figure compare 6 independently trained agents for both linear and multiplicative composition. Weights shown in Equation 12 are chosen so that undesirable performance on any criterion would have values near 0.
@@ VII-A Ablation Study on Multiplicative Composition
VIII Hyper-parameter search
We conducted a hyper-parameter search and a correlation factor analysis using 4 different prominent RL algorithms – PPO, DDPG, SAC and TD3 – for the quadrotor attitude control problem mainly considered in this paper. The list of tested hyper-parameters and their corresponding correlations to the training reward are presented in Tables IVVVI and VII respectively. Our hyper-parameter choices were informed by common values used to achieve the best performance across various benchmark Gymnasium environments [44].
Table III meanwhile compares the training performances of each of the algorithms, as well as DDPG×$\times$ which we introduce in this work, after 300000 training steps (corresponding to a the time by which DDPG×$\times$ typically starts to plateau), but on a slightly modified version of the problem – where agents are only required to minimize tracking error, without worrying about other characteristics such as control smoothness. This, we found, was typically a good starting point for tuning algorithms’ hyper-parameters and to determine if they are capable of addressing the underlying control problem. As the data demonstrates, all the tested algorithms perform significantly worse than DDPG×$\times$.
While the relative performance metrics appear similar across DDPG, PPO, SAC and TD3, there are critical differences in the quality of the behaviors being learned, which in turn impacts the ability of the algorithms to further improve agents’ performances. As shown in Fig. 11, the tested TF2 implementation of PPO does not behave similarly to the modified OpenAI Baselines [37] TF1 version used in previous work. [36, 35, 11]. Where the PPO based on Baselines learns to track the control signals well, albeit with high variance in the motor actuation (since smoothness is not being trained for), the TF2 implementation saturates the motor commands at the maximum, thus effectively doing nothing to control the drone’s attitude. This still generates some reward, since the training environment predominantly consists of smaller (but more frequently varying) control angular velocities. The problem with this, however, is that this behavior seems to be a local optima that is difficult for the algorithms to recover from to further improve performance.
When considering alternatives to PPO, we experimented with DDPG, SAC and TD3. As shown in Fig. 12, we observed that TD3 would often fall into similar trappings as PPO, with the motor outputs becoming saturated at maximum actuation. While DDPG was slower to learn (often performing worse than SAC), crucially, it did not saturate and would in fact consistently try to maintain control through more nuanced use of the motors – an extreme but illustrative case is shown in Fig. 12. Given our results, while SAC would have been preferable as a starting point based on the typical rewards achieved, the fact remained that training SAC was approximately 1.5 times slower than DDPG for relatively marginal gains, making it a less attractive option for iterating. We therefore chose DDPG to experiment on our proposed multiplicative policy optimization. In our tests, agents which achieve an average reward of more 0.8 were typically responsive enough to control in real flight.
TABLE IV: TF2 PPO paramater/reward correlations
@@ VIII Hyper-parameter search
TABLE V: DDPG paramater/reward correlations
@@
Figure 11: A snapshot of PPO and DDPG×$\times$ agents after 300,000 training steps acting in their training environment. Attitude control elements are split into Yaw, Pitch, and Roll in the left column. On the right are the outputs generated by the agents while attempting to follow the control inputs. Without smoothness considerations, within 300,000 steps, Baselines PPO performs very well on tracking while outputting actions with high variability (as is shown in previous works [36], [35]). Conversely, the TF2 implementation of PPO fails to learn useful behavior across hyper-parameters. DDPG×$\times$, which we introduced in this work, is able to reasonably follow the target while maintaining smooth outputs and continues to improve as training progresses.
TABLE VI: SAC paramater/reward correlations
@@
TABLE VII: TD3 paramater/reward correlations
@@
Figure 12: For the three DDPG-based algorithms, we observe clear dissimilarities in the quality of the policies learned, based on input tracking and motor usage. Entropy regularization employed in SAC allows trained agents to achieve relatively reasonable performance on both tracking and motor usage, but the algorithm was noted to not improve further with more training, likely due to the same regularization that initially helps preventing large responses, thus preventing the algorithm from learning to track more aggressive maneuvers. TD3 behaves similarly to the TF2 PPO, in that it tends to saturate at various local optima, outputting extreme values. DDPG, on the other hand, demonstrates some tracking capabilities can varies its control output drastically when close to the setpoint, similar in behavior to Baselines PPO. Given these three candidate algorithms and their relative benefits and drawbacks, we chose to modify DDPG as it runs faster than SAC and was capable of achieving good tracking performance. While DDPG itself was not consistently reliable, DDPG×$\times$, with our proposed improvements, was repeatable and controllable.
@@ IX Known Limitations
While anchors clearly reduce power consumption without resulting in catastrophic forgetting, the adapted agents were not observed to improve their tracking performance significantly, even though this was one of the key learning objectives of the policy optimization. It is as yet unclear as to whether this is simply a limitation of the DDPG algorithm (or our specific implementation), or a matter of how the regularization was tuned. In cases where domain gaps are large, the very stability offered by the anchor critics may significantly compromise fine-tuning of behavior and could result in policies that are ill-suited for both the initial and transferred domain. Finally, our usage of a saved replay buffer from simulation to evaluate anchor critics limits the ability to adapt anchors to new policies, as a divergent enough policy will make anchors an unsuitable q-value for the source domain. We suspect that keeping the simulation in the loop might allow for anchors to allow for continued learning long-term.

### v3

@@ IV-B Catastrophic Forgetting in Reacher
This issue is especially grave on the reacher task. Here, the target domain has identical dynamics to the source, differing only in its goal distribution: target setpoints are restricted to a strict subset of the source domain’s broader range (Table I). This creates a limited, skewed task distribution during fine-tuning. The resulting instability of the naively fine-tuned agent (Fig. 4) stems from overfitting to this limited target distribution and consequently forgetting how to handle the wider range of goals present in the source domain.
@@ V Anchor Critics in Reality
    Leveraging SwaNNL, we evaluate how live-adapting with and without anchors affects safety, smoothness, and performance when adapting policies from simulation to a real drone system (Sections VI-B and VI-C). We focus on how anchors mitigate issues arising from adapting to potentially limited real-world data distributions. The novel infrastructure offered by SwaNNL enabled us to develop and test practical ways to adapt agents while a quadrotor is in flight, demonstrating how anchors make live adaptation safer and more robust against forgetting intended behaviors.
@@ V-A Engineering SwaNN Lake for Live NN updates
In this work, we investigate strategies for RL-based live control adaptation on real robots. As such, two important sub-goals are formulated as follows. First, we needed to allow BetaFlight-compatible firmware to update its NN controller(s) without interrupting the control loop. As BetaFlight is a standard across various robotic platforms, this allows our framework to support a large range of robots, such as blimps, boats, helicopters, racing drones, and more. Second, we needed to be able adapt the agents live without crashing. This necessitated preventing catastrophic forgetting, particularly given the risk of overfitting to potentially skewed or limited data distributions encountered during live operation.
We thus decided to evolve Neuroflight [10, 11], initially developed by Koch et al. for allowing the use of NNs for high-performance quadrotor control. Neuroflight extends the open-source Betaflight flight-control firmware stack allowing NN models to be compiled and embedded into the control firmware. We evolved the Neuroflight framework such that modifications to the employed NN controller(s) can be carried out without interrupting the control loop. To do so, we re-designed the firmware so that we could swap out binarized NN graphs with minimal overhead rather than having them directly compiled into the firmware. In particular, it is possible to both (i) update the NN controller weights and (ii) swap-in graphs with different NN architectures. Our novel firmware stack outlined in Fig. 6, is open-sourced [12]. Our architecture allows inference to be performed onboard the micro-controller, but training onboard is prohibitively expensive, given compute and storage limitations of low-power embedded platforms. We thus introduce a point-to-point communication channel with an external ground station, to mitigate this issue. Most importantly, the ground station can be used as a training platform to adapt controllers based on data captured in flight. SwaNNFlight further integrates with wireless transceiver modules to (i) wirelessly transmit flight data and (ii) accept new NN models to switch to mid-flight.
@@ VI TRAINING FLIGHT CONTROLLERS WITH SwaNNL
Transitioning to TF2 introduced a problem for the training pipeline. Prior works [19, 18, 11] train flight-control agents with the OpenAI Baselines implementation of PPO [9]. While reproducible with the software packages and versions reported by the authors, the performance observed with PPO agents in TF1 was not achievable by any of several tested TF2 implementations of PPO, including TF-agents, stable-baselines etc., despite using identical network architectures and performing a full hyperparameter search. Algorithms like PPO and TRPO are known to be sensitive to (and even brittle under) implementation specifics [20], prompting a search for alternatives.
Viable flight-control policies were trainable using a TF2 translation of OpenAI’s Spinning Up PyTorch implementation of DDPG [15]. While vanilla DDPG presented with similar oscillatory behavior identified in the work on CAPS  [19], introducing CAPS regularization to DDPG enabled the training of viable agents in simulation. We found, however, that the performance of DDPG on the control problem at hand exhibited relatively high variance: not all the trained agents were safely flyable. This prompted us to explore alternative formulations which improve upon CAPS defined in the next section.
@@ VI-B Catastrophic Forgetting and Safety Concerns in Adaptation without Anchors
Before testing adaptation in live, unconstrained flight, we first conducted a series of tests in a controlled lab environment where the drone was attached to a light tether to restrain it in the event of a catastrophic loss of control, without being a hindrance to flight. These tests consisted of fine-tuning a well-trained controller from simulation (the source domain) on the real platform (the target domain) while giving it typically small control targets ($<50$ deg/s) and occasionally requesting control in excess of that ($>100$ deg/s). Without anchors preserving the agent’s performance profile across the full range of control inputs defined in simulation, the live agents were observed to forget how to maintain tracking when receiving large inputs, as this occurs less frequently during adaptation. Such large inputs are common during acrobatic maneuvers, which are common in simulation tasks but not in real-world flights. We would observe agents sometimes regaining control, and often losing it again, as seen in Fig. 7. More importantly, this also results in a sudden loss of control where errors would quickly snowball and motor actuation rises sharply, risking damage to the drone. Fig. 8 demonstrates how, when commanding attitudes not seen for a while during live-flight adaptation, agents could very quickly lose control (observed from MAEs increasing quickly and exponentially). We attributed this to overfitting to the limited or skewed distribution of the live-flight training data (predominantly small control inputs), similar to the Reacher environment. Anchors, by preserving the intended response learned across the diverse simulation distribution, help in avoiding this class of problems.
@@ VI-C Live Adaptation During Unconstrained Flight
We use two key performance metrics in policy evaluation. (1) Tracking accuracy is measured using the Mean Absolute Error (MAE). (2) Smoothness is computed by taking the Fast Fourier Transform (FFT) of the actuation commands, and then computing $Sm$ as defined by Equation 4 in [19] (lower is better). Shown in Fig. 10 is a comparison of control signals before and after 15 adaptation steps with anchor critics. Motor usage is reduced significantly, resulting in improved control smoothness and almost a 50% reduction in power consumption. Crucially, this is achieved with comparable MAEs to the original simulated agent (Table II), indicating that tracking performance is maintained across the intended operational range, thanks to the anchor preserving the behavior learned on the diverse simulation distribution. Table II further details how MAE, smoothness, and current draw change through adaptation.
VII Applying Multi-Objective Composition Principles from FPL
Linearly combining multiple RL objectives often obscures designer intent. Fulfillment Priority Logic (FPL) [14] was introduced to provide principled composition operators for consolidating multi-objective optimization in RL, with one instance being the use of the geometric mean $\boldsymbol{\land}^{0}$ as a smooth AND operator. The FPL $\boldsymbol{\land}^{0}$ operator explicitly targets the simultaneous satisfaction of objectives, represented as objectives with fulfillment levels between $[0,1]$. By making use of this framework, we can scalarize our multi-objective policy optimization as a single objective optimization problem. Since we propose to simulatenously satisfy multiple objectives over both the simulation and real-world domains, we can use the operator provided by FPL to compose our objective functions for both domains into a single scalarized objective function.
We take the standard expected task fulfillment (normalized Q-value, $Q_{\pi_{\theta}}$) and the temporal ($L_{T}$) and spatial ($L_{S}$) smoothness penalties from CAPS [19], along with an actuation penalty term derived from the pre-activation output ($\pi_{\theta}^{:-1}$). These penalty terms are normalized into fulfillment values $F_{T},F_{S},F_{A}\in[0,1]$, where higher values represent better fulfillment (or satisfaction of our objectives). The policy optimization then maximizes the FPL conjunction of these fulfillments:
$J_{\pi_{\theta}}=\boldsymbol{\land}^{0}(Q_{\pi_{\theta}}(s,a),F_{T},F_{S},F_{A}).$ (10)
@@ VII Applying Multi-Objective Composition Principles from FPL
As demonstrated empirically (Section VII-A, Fig. 9), this FPL-based composition significantly reduces training variance compared to linear approaches.
VII-A Ablation Study on Applying FPL-based Composition
To validate the effectiveness of applying FPL-based multiplicative composition shown in Equation 10 in our specific DDPG setup, we trained controllers in simulation comparing the linear loss composition from CAPS [19] against FPL-based multiplicative composition. We conducted extensive hyperparameter search with different weighting parameters, both for the CAPS regularization terms as well as the training rewards. We show in Fig. 9 that we achieved higher performance with less tuning and significantly reduced variance using multiplicative composition. The results presented in the figure compare 6 independently trained agents for both linear and multiplicative composition.
@@ VII-A Ablation Study on Applying FPL-based Composition
TABLE IV: TF2 PPO parameter/reward correlations
@@
TABLE V: DDPG parameter/reward correlations
@@
TABLE VI: SAC parameter/reward correlations
@@
TABLE VII: TD3 parameter/reward correlations
@@
VIII Hyper-parameter search
[…]
@@ IX Known Limitations
(1) In cases where domain gaps are large, the very stability offered by the anchor critics may significantly compromise fine-tuning of behavior and could result in policies that are ill-suited for both the initial and transferred domain.
(2) While anchors clearly reduce power consumption without resulting in catastrophic forgetting, the adapted agents were not observed to improve their tracking performance significantly, even though this was one of the key learning objectives of the policy optimization. We hypothesize that this is due to the fact that we relied on replay buffers from simulation to evaluate anchor critics, rather than have a continuously updated anchor via a simulation in the loop, we deem this a promising direction for future work.
(3) The complexity of implemention may detract from the wide-applicability of the approach, even though we provide a generic implementation applicable over a range of robotic platforms via SwaNNFlight and SwaNNL, orchestrating data from simulation and real-world data is non-trivial.
X Acknowledgements
We thank Patrick Carpanedo for his contributions to the video presentation and adding polishing touches to the manuscript.